# Effectively Leveraging Momentum Terms in Stochastic Line Search Frameworks for Fast Optimization of Finite-Sum Problems

### v1

@@ 4 The Algorithmic Framework
17:     Set $\beta_{k+1}=\frac{\|\nabla f_{\mathcal{S}_{k}}(x^{k+1})\|^{2}}{\|\nabla f_{\mathcal{S}_{k}}(x^{k})\|^{2}}$
18:end for
4.1 Convergence aspects
At this point, it would be nice to provide some convergence result concerning Algorithm 1. Unfortunately, in this section we will only be able to highlight the difficulties associated with such an analysis.
The main issue stems from the employment of the data persistency strategy. The overlap between consecutive mini-batches allows to make functions $f_{k}$ and $f_{k-1}$ more similar to each other; in formal terms, what we are actually doing is to reduce the variance of $f_{k}$, conditioned to $x^{k}$, at the cost of the introduction of a bias. Indeed, assuming for simplicity $|\mathcal{S}_{k}|=|\mathcal{S}_{k-1}|$, we have
$\mathbb{E}\left[\nabla f_{k}(x)\mid x^{k}\right]=\mathbb{E}\left[\frac{1}{2}\nabla f_{\mathcal{S}_{k}}(x)+\frac{1}{2}\nabla f_{\mathcal{S}_{k-1}}(x)\mid x^{k}\right]\\ =\frac{1}{2}\mathbb{E}\left[\nabla f_{\mathcal{S}_{k}}(x)\mid x^{k}\right]+\frac{1}{2}\mathbb{E}\left[\nabla f_{\mathcal{S}_{k-1}}(x)\mid x^{k}\right]=\frac{1}{2}\nabla f(x)+\frac{1}{2}\nabla f_{\mathcal{S}_{k-1}}(x).$
This fact immediately breaks one of the most fundamentals assumptions typically made for the analysis of SGD type methods (see, e.g., [4]). The convergence analysis with biased gradient estimators would thus go far beyond the scope of this work.
If, on the other hand, we gave up on the the data persistency strategy, it would be possible to guarantee fast convergence of the algorithmic framework under interpolation and PL assumptions. To this aim, we would need to strengthen the safeguard condition on the direction. Instead of checking that $d$ is of descent for $f_{k}$, i.e., $d^{T}\nabla f_{k}(x^{k})<0$, we can instead check that the direction is stochastic gradient related, i.e.,
$\|d\|\leq c_{1}\|\nabla f_{k}(x^{k})\|\qquad\text{and}\qquad d^{T}\nabla f_{k}(x^{k})\leq-c_{2}\|\nabla f_{k}(x^{k})\|^{2},$
for some given constants $c_{1},c_{2}>0$. This property can be enforced, when not satisfied by the CG type direction, switching to the stochastic gradient direction or by means of the subspace minimization strategy from [24]. Then the convergence result would follow from the analysis in [35].
@@ 5 Computational Experiments
In this section, we present the results of a comprehensive experimental analysis of the proposed Mini-Batch Conjugate Gradient method with Data Persistency (MBCG-DP). We first compare its possible configurations and then evaluate its performance in comparison to state-of-the-art approaches. All experiments reported in this section were implemented in Python3333The code for the proposed method is available at github.com/dadoPuccio/MB-Conjugate-Gradient-DP and run on a machine with the following characteristics: Ubuntu 22.04, Intel i5-13400F Processor, 32 GB RAM, Nvidia GeForce RTX 4060 8GB.
@@
In Figure LABEL:fig:CG_rule_comparison we reported the behavior of MBCG-DP when comparing the CG rules of Fletcher-Reeves (FR), Polyak-Polak-Ribiére (PPR) and Hestenes-Stiefel (HS). In this experiment the generalized SPS (7) with $c=1$ is used to select $\alpha_{0}^{k}$ and the momentum clipping approach (contraction factor 0.5) from [20] is employed in case $d_{k}$ is not a descent direction. The Fletcher-Reeves rule clearly resulted the most effective one in both tasks and for both batch sizes.
@@
In Figure LABEL:fig:CG_tentative_step_comparison we show the behavior of MBCG-DP when using different approaches for selecting the tentative step $\alpha_{0}^{k}$ for the line search: a constant value, the heuristic employed in [17, 20] and the generalized SPS (7) with $c=1$ were tested. The Fletcher-Reeves rule is used in this experiment; in case $d_{k}$ is not a descent direction the repeated momentum clipping approach is used, again with contraction factor of $0.5$. The generalized SPS proved to be the most effective choice.
Finally, in Figure LABEL:fig:CG_recover_direction we observe the impact of different recovery options to retrieve a descent direction, namely: switching to the simple (negative) stochastic gradient direction, recomputing the direction according to the subspace optimization strategy proposed in [24], inverting the direction and taking $-d_{k}$ and adopting the repeated momentum clipping approach (contraction factor of $0.5$) from [20]. In this experiment, the Fletcher-Reeves rule is employed in combination with the generalized SPS initial step size. The clipping strategy resulted the preferable choice especially in nonconvex problems.
This detailed comparison allows us to identify a promising setup to run MBCG-DP in further experiments to compare it to state-of-the-art methods. In what follows, we thus employed the Fletcher-Reeves rule for $\beta_{k}$, the generalized stochastic Polyak step size (7) for choosing $\alpha_{0}^{k}$ and the clipping strategy on $\beta_{k}$ as a recovery option for obtaining a descent direction.
The other algorithms included in the following experiments are: vanilla SGD with momentum, Adam [8], SLS [17], PoNoS [18] and MSL_SGDM [20]. The benchmark of test problems include three convex instances, consisting in the training of an RBF-kernel classifier on mushrooms, ijcnn and rcv1 datasets444datasets available at csie.ntu.edu.tw/ cjlin/libsvmtools/datasets. Moreover, three nonconvex learning tasks are taken into account: once again a multi-layer perceptron with 1000 hidden units and ReLU activations to be trained on MNIST, a three-layered convolutional architecture with 32 channels each, followed by a fully-connected layer of 512 units to be trained on FashionMNIST [36] dataset555available at github.com/zalandoresearch/fashion-mnist and a ResNet18 architecture to be trained on CIFAR10 [37] dataset666available at cs.toronto.edu/ kriz/cifar.html. For the latter network, we employed the normalization-free setup proposed in [38]. Batch sizes of 128 and 512 were again considered for all problems.
Table 2: Configurations in which a 50% overlap resulted in a better performing algorithm.
@@
SGD+M
Adam
SLS
PoNoS
MSL_SGDM
@@
Figure 4: Comparison of the proposed MBCG-DP algorithm with other algorithms in the literature. The training loss over time is shown for each problem.
@@
In order to offer a fair comparison between the proposed MBCG-DP method and its competitors, each reference algorithm was evaluated to assess whether a 50% overlap between mini-batches could enhance its performance for each problem considered. An algorithm is deemed to perform better with the overlap if it either terminates before its non-overlapped counterpart with a superior solution or, at the termination time of the non-overlapped algorithm, the overlapped one holds a better solution. In both cases the training loss is compared to asses whether a solution is better than the other. We report in Table 2 the outcome of this preliminary analysis. In convex problems, we observe that overlap is almost always beneficial. However, with larger architectures in nonconvex tasks, this benefit does not appear to hold. We argue that, in such cases, the dominant computational burden shifts from data I/O operations to the evaluation of the training loss and its gradient. Since the number of iterations per epoch with a 50% overlap is doubled, in this cases the execution time per epoch will also double. In addition, we remark that good hyperparameters for the standard algorithm may need adjustment to enhance performance when overlap is employed.
@@
Figure 5: Comparison of the proposed MBCG-DP algorithm with other algorithms in the literature. The validation accuracy over time is shown for each problem.

### v3

@@ 5 The Algorithmic Framework
17:  Select $\mathcal{R}_{k}\subset B_{k}$ according to a deterministic rule
18:  Define $\beta_{k+1}$ from $\nabla f_{\mathcal{R}_{k}}(x^{k})$ and $\nabla f_{\mathcal{R}_{k}}(x^{k+1})$ according to any CG rule (11)
19:end for
5.1 Convergence Analysis
We now turn to the presentation of a formal convergence analysis for Algorithm 1. The analysis tackles two main issues, that are addressed in two separate parts.
5.1.1 Estimator Bias in Presence of Mini-Batch Persistency
The key element of most convergence analyses for stochastic methods, including the one we carry out in the next subsection, relies on the availability of a conditionally unbiased estimator $g_{k}(x)$ of $\nabla f(x^{k})$, i.e. $\mathbb{E}_{k}[g_{k}(x^{k})]=\nabla f(x^{k})$. We use $\mathbb{E}_{k}[\cdot]$ to denote the conditional expectation with respect to $x^{k}$, meaning $\mathbb{E}_{k}[\cdot]=\mathbb{E}[\cdot\mid x^{k}]$, while $\mathbb{E}[\cdot]$ indicates the total expectation.
However, in contrast to classical setups, where the unbiased estimator is reasonably obtained by drawing at each iteration the mini-batch to be used completely at random, we employ a data persistency strategy that turns this assumption false, as the set of indexes $\mathcal{R}_{k-1}$ is fixed, and thus deterministic, for the iteration $k$.
Let us denote $\bar{\mathcal{S}}=\{1,\ldots,N\}\setminus\mathcal{S}$. Recalling that $\nabla f_{\mathcal{S}}(x)=\frac{1}{|\mathcal{S}|}\sum_{i\in{\mathcal{S}}}\nabla f_{i}(x)$, and assuming for simplicity that $\mathcal{S}_{k}$ is randomly drawn from $\bar{\mathcal{R}}_{k-1}$ - so that $\mathbb{E}_{k}[\nabla f_{\mathcal{S}_{k}}(x^{k})]=\nabla f_{\bar{\mathcal{R}}_{k-1}}(x^{k})$ - we get that
$\mathbb{E}_{k}[g_{k}(x^{k})]$ $=\mathbb{E}_{k}\left[\frac{1}{|B_{k}|}\sum_{i\in{B_{k}}}\nabla f_{i}(x^{k})\right]$ (12)
$=\mathbb{E}_{k}\left[\frac{1}{|B_{k}|}\left(\sum_{i\in{\mathcal{R}}_{k-1}}\nabla f_{i}(x^{k})+\sum_{i\in{\mathcal{S}}_{k}}\nabla f_{i}(x^{k})\right)\right]$
$=\frac{1}{|B_{k}|}\sum_{i\in{\mathcal{R}}_{k-1}}\nabla f_{i}(x^{k})+\mathbb{E}_{k}\left[\frac{1}{|B_{k}|}\sum_{i\in{\mathcal{S}}_{k}}\nabla f_{i}(x^{k})\right]$
$=\frac{|\mathcal{R}_{k-1}|}{|B_{k}|}\nabla f_{\mathcal{R}_{k-1}}(x^{k})+\frac{|\mathcal{S}_{k}|}{|B_{k}|}\mathbb{E}_{k}\left[\nabla f_{\mathcal{S}_{k}}(x^{k})\right]$
$=\frac{|\mathcal{R}_{k-1}|}{|B_{k}|}\nabla f_{\mathcal{R}_{k-1}}(x^{k})+\frac{|\mathcal{S}_{k}|}{|B_{k}|}\nabla f_{\bar{\mathcal{R}}_{k-1}}(x^{k}).$
If, for instance, we assume $|\mathcal{R}_{k-1}|=|\mathcal{S}_{k}|=|B_{k}|/2$, we get
$\mathbb{E}_{k}[g_{k}(x^{k})]=\frac{1}{2}\nabla f_{\mathcal{R}_{k-1}}(x^{k})+\frac{1}{2}\nabla f_{\bar{\mathcal{R}}_{k-1}}(x^{k}),$
which is a gradient estimate that is clearly biased towards the “persistent” samples. When taking the expected value for the gradient estimator, the weight assigned to the terms $\nabla f_{i}(x^{k})$ for $i\in\mathcal{R}_{k-1}$ is $\frac{1}{2|\mathcal{R}_{k-1}|}$, evidently higher than it would be in the common unbiased stochastic gradient, i.e., $\frac{1}{N}$.
Fortunately, we can recover unbiasedness of the estimator applying a suitable correction when constructing the approximation $g_{k}(x^{k})$ - and similarly $f_{k}(x^{k})$. This result is formalized hereafter.
Proposition 5.1.
Let $\mathcal{R}_{k-1}\subset\{1,\ldots,N\}$ and let $\mathcal{S}_{k}$ be a random subset of $\bar{\mathcal{R}}_{k-1}=\{1,\ldots,N\}\setminus\mathcal{R}_{k-1}$. Let $\zeta_{k}=\frac{N-|\mathcal{R}_{k-1}|}{|\mathcal{S}_{k}|}$. Then, the quantities
$f_{k}(x^{k})=\frac{1}{N}\left(\sum_{i\in{\mathcal{R}}_{k-1}}f_{i}(x^{k})+\zeta_{k}\sum_{i\in{\mathcal{S}_{k}}}f_{i}(x^{k})\right)$
and
$g_{k}(x^{k})=\frac{1}{N}\left(\sum_{i\in{\mathcal{R}}_{k-1}}\nabla f_{i}(x^{k})+\zeta_{k}\sum_{i\in{\mathcal{S}_{k}}}\nabla f_{i}(x^{k})\right)$
are conditionally unbiased estimators of $f(x^{k})$ and $\nabla f(x^{k})$ given $x^{k}$, i.e., $\mathbb{E}_{k}[f_{k}(x^{k})]=f(x^{k})$ and $\mathbb{E}_{k}[g_{k}(x^{k})]=\nabla f(x^{k})$.
Proof.
We only prove the result for $g_{k}$, as the proof for $f_{k}$ is substantially identical. Taking the conditional expected value of $g_{k}(x^{k})$ given $x^{k}$, recalling that $\frac{1}{|\mathcal{S}|}\sum_{i\in{\mathcal{S}}}\nabla f_{i}(x)=\nabla f_{\mathcal{S}}(x)$, that $\mathbb{E}_{k}[\nabla f_{\mathcal{S}_{k}}(x^{k})]=\nabla f_{\bar{\mathcal{R}}_{k-1}}(x^{k})$ and (12), we get
$\mathbb{E}_{k}[g_{k}(x^{k})]$ $=\mathbb{E}_{k}\left[\frac{1}{N}\left(\sum_{i\in{\mathcal{R}}_{k-1}}\nabla f_{i}(x^{k})+\zeta_{k}\sum_{i\in{\mathcal{S}}_{k}}\nabla f_{i}(x^{k})\right)\right]$
$=\frac{|\mathcal{R}_{k-1}|}{N}\nabla f_{\mathcal{R}_{k-1}}(x^{k})+\zeta_{k}\frac{|\mathcal{S}_{k}|}{N}\nabla f_{\bar{\mathcal{R}}_{k-1}}(x^{k})$
$=\frac{1}{N}\left(\sum_{i\in{\mathcal{R}}_{k-1}}\nabla f_{i}(x^{k})+\frac{N-|\mathcal{R}_{k-1}|}{|\mathcal{S}_{k}|}\frac{|\mathcal{S}_{k}|}{N-|\mathcal{R}_{k-1}|}\sum_{i\in{\bar{\mathcal{R}}}_{k-1}}\nabla f_{i}(x^{k})\right)$
$=\frac{1}{N}\sum_{i=1}^{N}\nabla f_{i}(x^{k})=\nabla f(x^{k}),$
thus getting the thesis. ∎
Remark 2.
The correction to make the gradient estimate unbiased even in presence of mini-batch persistency allows to soundly employ the latter technique from a theoretical convergence standpoint. However, we note that the correction heavily alleviates the sought effects that drove the introduction of the overlap strategy itself in the first place, since the weight assigned to the “persistent” samples is drastically lowered. This is also evident in the computational experiments, as shown later in Figure 3. Therefore, we believe the theoretical analysis of convergence with a biased estimator would also be of great interest. Yet, this issue is a large open problem and its treatment would go far beyond the scope of this work; we thus leave it to future studies.
5.1.2 Properties of the Search Direction
Together with the assumption on gradient estimate unbiasedness, largely discussed in the previous subsection, we need to state another important assumption for properly analyzing the algorithm: the search directions $d_{k}$ shall in fact be somehow tied to the true gradient $\nabla f(x^{k})$. As pointed out in well-established literature (see, e.g., [bottou2018, sec. 4.1]), the sequence of search directions $\{d_{k}\}$ shall be such that, for all $k$,
$\mathbb{E}_{k}[\|d_{k}\|]\leq C_{1}\|\nabla f(x^{k})\|,\quad\text{ and }\quad\mathbb{E}_{k}[d_{k}]^{T}\nabla f(x^{k})\leq-\|\nabla f(x^{k})\|^{2}.$
Following the discussion in [lapucci2024convergenceconditionsstochasticline], this can be ensured if, given constants $c_{1}$, $c_{2}$ and $c_{3}$, the entire sequence $\{d_{k}\}$ satisfies the following set of conditions:
$\|d_{k}\|\leq c_{1}\|g_{k}(x^{k})\|,$ (13)
$d_{k}^{T}g_{k}(x^{k})\leq-c_{2}\|g_{k}(x^{k})\|^{2},$ (14)
$\operatorname{Var}_{k}(d^{k})\leq c_{3}\operatorname{Var}_{k}(g_{k}(x^{k})),$ (15)
where by $\operatorname{Var}_{k}$ we denote the conditional variance of a vector, i.e., $\operatorname{Var}_{k}(v)=\mathbb{E}_{k}[\|v\|^{2}]-\|\mathbb{E}_{k}[v]\|^{2}$.
Properties (13)-(14) can be checked, as safeguards, at each iteration for fixed values of $c_{1}$ and $c_{2}$: when not satisfied by the CG type direction, the algorithm can switch to the stochastic gradient direction.
In order to ensure that condition (15) is also satisfied by the direction resulting from the safeguard condition, we need to keep parameter $\beta_{k}$ bounded; this is also not particularly restrictive, as it can be enforced by a clipping operation at the update time; in other words, we can simply set $\beta_{k+1}=\min\{\hat{\beta}_{k+1},\bar{\beta}\}$, where $\hat{\beta}_{k+1}$ is the value resulting from the employed CG update rule. With this precaution, we only have to rely on an assumption on the variance of $\beta_{k+1}$, which appears reasonable as argued in [lapucci2024convergenceconditionsstochasticline, Remark 2].
The satisfaction of condition (15) for the whole sequence, if clipping and safeguards are suitably applied, is formally stated and proved in the next proposition.
Proposition 5.2.
Let, for all $k$, $d_{k}=-g_{k}(x^{k})+\beta_{k}d_{k-1}$, where
$\beta_{k}=\begin{cases}\min\{\hat{\beta}_{k},\bar{\beta}\}&\text{if }\|p_{k}\|\leq c_{1}\|g_{k}(x^{k})\|\text{ and }p_{k}^{T}g_{k}(x^{k})\leq-c_{2}\|g_{k}(x^{k})\|^{2},\\ 0&\text{otherwise},\end{cases}$
with $p_{k}=-g_{k}(x^{k})+\min\{\hat{\beta}_{k},\bar{\beta}\}d_{k-1}$. Assuming that, for all $k$, $\|d_{k-1}\|\leq M$ and $\operatorname{Var}_{k}(\beta_{k})\leq\hat{c}_{3}\operatorname{Var}_{k}(g_{k}(x^{k}))$ for some constants $M,\hat{c}_{3}>0$, we have that $\operatorname{Var}_{k}(d^{k})\leq c_{3}\operatorname{Var}_{k}(g_{k}(x^{k}))$, with $c_{3}=\left(1+M^{2}\hat{c}_{3}+2M\sqrt{\hat{c}_{3}}\right)$.
Proof.
The variance of the direction $d_{k}$ is given by
$\operatorname{Var}_{k}(-g_{k}(x^{k})+\beta_{k}d_{k-1})=\operatorname{Var}_{k}(g_{k}(x^{k}))+\|d_{k-1}\|^{2}\operatorname{Var}_{k}(\beta_{k})-2\operatorname{Cov}_{k}(g_{k}(x^{k}),\beta_{k}d_{k-1})$
Exploiting the Cauchy-Schwartz inequality on the dot product $u^{T}v$ where $u_{i}=\sqrt{\operatorname{Var}_{k}((g_{k}(x^{k}))_{i})}$ and $v_{i}=|(d_{k-1})_{i}|$, we can write
$|\operatorname{Cov}_{k}(g_{k}(x^{k}),\beta_{k}d_{k-1})|$ $\leq\sum_{i=1}^{n}\left|\operatorname{Cov}_{k}((g_{k}(x^{k}))_{i},\beta_{k}(d_{k-1})_{i})\right|$
$\leq\sum_{i=1}^{n}\sqrt{\operatorname{Var}_{k}((g_{k}(x^{k}))_{i})}\sqrt{\operatorname{Var}_{k}(\beta_{k}(d_{k-1})_{i})}$
$=\sqrt{\operatorname{Var}_{k}(\beta_{k})}\sum_{i=1}^{n}\sqrt{\operatorname{Var}_{k}((g_{k}(x^{k}))_{i})}|(d_{k-1})_{i}|$
$\leq\sqrt{\operatorname{Var}_{k}(\beta_{k})}\sqrt{\operatorname{Var}_{k}(g_{k}(x^{k}))}\|d_{k-1}\|.$
Using that $\|d_{k-1}\|\leq M$ and that $\operatorname{Var}_{k}(\beta_{k})\leq\hat{c}_{3}\operatorname{Var}_{k}(g_{k}(x^{k}))$ we obtain
$\operatorname{Var}_{k}(d_{k})$ $\leq\operatorname{Var}_{k}(g_{k}(x^{k}))+\|d_{k-1}\|^{2}\operatorname{Var}_{k}(\beta_{k})+2\sqrt{\operatorname{Var}_{k}(\beta_{k})}\sqrt{\operatorname{Var}_{k}(g_{k}(x^{k}))}\|d_{k-1}\|$
$\leq\left(1+M^{2}\hat{c}_{3}+2M\sqrt{\hat{c}_{3}}\right)\operatorname{Var}_{k}(g_{k}(x^{k})),$
which concludes the proof. ∎
Proposition 5.2 provides the final building block for establishing convergence results for the proposed algorithm. Note that the first assumption on $\|d_{k-1}\|$ is not restrictive, as it can be enforced at iteration $k$ by simple clipping of the then deterministic direction $d_{k-1}$ (i.e., using $\min\{1,\frac{M}{\|d_{k-1}\|}\}d_{k-1}$ instead of $d_{k-1}$).
5.1.3 Convergence Results
We are now ready to present the asymptotic convergence analysis for Algorithm 1. In particular, we are showing how the method proposed in this paper possesses analogous properties as the other well-known stochastic line-search based methods from the literature [vaswani2019fast, vaswani2019, loizou2021, galli2023, lapucci2024convergenceconditionsstochasticline]. Specifically, it is possible to guarantee fast convergence of the algorithmic framework under interpolation and PL assumptions [ma2018, mishkin2020interpolation, polyak1987introduction].
We recall that the minimizer interpolation property [mishkin2020interpolation, Def. 3] states that
$x^{*}\in\operatorname*{\arg\min}_{x^{*}\in\mathbb{R}^{n}}f(x)\implies x^{*}\in\operatorname*{\arg\min}_{x^{*}\in\mathbb{R}^{n}}f_{i}(x)\text{ for all }i\in\{1,\dots,N\}.$
On the other hand, we say that a function $f$ satisfies the PL condition if
[…]
We shall also recall that (see, e.g., [lapucci2024convergenceconditionsstochasticline, Sec. 3]), $L$-smoothness, minimizers interpolation and the PL condition imply the Strong Growth Condition to hold for some $\rho>0$, i.e.,
$\mathbb{E}_{k}[\|\nabla f_{k}(x)\|^{2}]\leq\rho\|\nabla f(x)\|^{2}\text{ for all }x\in\mathbb{R}^{n}\text{ and for any }k.$
Since we need to insert into the algorithmic framework the fixes described in sections 5.1.1 and 5.1.2, we report for the sake of clarity the instructions of the provably convergent version of the algorithm in Algorithm 2. Note that we also made explicit the constant lower bound required for the initial step size $\alpha_{k}^{0}\geq\bar{\alpha}$ of the line search procedure.
Algorithm 2 Convergent MBCG-DP
1:Input: $x^{0}\in\mathbb{R}^{n}$, $f_{1},\dots,f_{N}:\mathbb{R}^{n}\rightarrow\mathbb{R}$, $\gamma\in(0,1)$, $\delta\in(0,1)$, $c_{1},c_{2},\bar{\alpha},\bar{\beta},M>0$.
2:Let $\mathcal{R}_{0}=\emptyset$,
[…]
4:for $k=1,2,\ldots$ do
5:  Randomly draw a subset $\mathcal{S}_{k}\subset\{1,\ldots,N\}\setminus\mathcal{R}_{k-1}$
6:  Set $B_{k}=\mathcal{R}_{k-1}\cup\mathcal{S}_{k}$
7:  Set $\zeta_{k}=\frac{N-|\mathcal{R}_{k-1}|}{|\mathcal{S}_{k}|}$
8:  Let $f_{k}=\frac{1}{N}\left(\sum_{i\in{\mathcal{R}}_{k-1}}f_{i}+\zeta_{k}\sum_{i\in{\mathcal{S}_{k}}}f_{i}\right)$
9:  Let $g_{k}=\frac{1}{N}\left(\sum_{i\in{\mathcal{R}}_{k-1}}\nabla f_{i}+\zeta_{k}\sum_{i\in{\mathcal{S}_{k}}}\nabla f_{i}\right)$
10:  Set $d=-{g}_{k}(x^{k})+\beta_{k}\min\left\{1,\frac{M}{\|d_{k-1}\|}\right\}d_{k-1}$
11:  if $\|d\|\leq c_{1}\|g_{k}(x^{k})\|$ and $d^{T}g_{k}(x^{k})\leq-c_{2}\|g_{k}(x^{k})\|^{2}$ then
12:   Set $d_{k}=d$
13:  else
14:   Set $d_{k}=-g_{k}(x^{k})$
15:  end if
16:  Define an initial step size $\alpha_{0}^{k}\geq\bar{\alpha}$
17:  Set $\alpha_{k}=\max_{j=0,1,\ldots}\{\alpha_{0}^{k}\delta^{j}\mid f_{k}(x_{k}+\alpha_{0}^{k}\delta^{j}d_{k})\leq f_{k}(x_{k})+\gamma\alpha_{0}^{k}\delta^{j}g_{k}(x_{k})^{T}d_{k}\}$
18:  Set $x_{k+1}=x_{k}+\alpha_{k}d_{k}$
19:  Select $\mathcal{R}_{k}\subset B_{k}$ according to a deterministic rule
20:  Define $\hat{\beta}_{k+1}$ from $\nabla f_{\mathcal{R}_{k}}(x^{k})$ and $\nabla f_{\mathcal{R}_{k}}(x^{k+1})$ according to any CG rule (11)
21:  Set $\beta_{k+1}=\min\{\hat{\beta}_{k+1},\bar{\beta}\}$
22:end for
The main convergence theorem is stated hereafter. The result is recovered by [lapucci2024convergenceconditionsstochasticline, Th. 5.3].
Theorem 5.3.
Assume $f:\mathbb{R}^{n}\to\mathbb{R}$ is an $L$-smooth function, satisfying the PL condition with constant $\mu$ and the minimizer interpolation property, and that the randomly drawn functions $f_{k}$ considered at any iteration $k$ are $L_{k}$-smooth functions, with $L\leq L_{\text{max}}=\max_{k}L_{k}$. Let the sequence $\{\beta_{k}\}$ be such that $\operatorname{Var}_{k}(\beta_{k})\leq\hat{c}_{3}\operatorname{Var}_{k}(g_{k}(x^{k}))$ and that $c_{2}>c_{3}\left(1-\frac{1}{\rho}\right)$, being $c_{3}=\left(1+M^{2}\hat{c}_{3}+2M\sqrt{\hat{c}_{3}}\right)$ and $\rho$ the SGC constant. Then, if $\{x^{k}\}$ is the sequence produced by Algorithm 2 the following property holds:
$\mathbb{E}[f(x^{k+1})-f(x^{*})]\leq(\eta\bar{\alpha})^{k}(f(x^{0})-f(x^{*})),$ (16)
where $\eta=\left(\frac{L_{\text{max}}c_{1}^{2}}{2c_{2}}\left(\frac{1}{\gamma}+\frac{1}{\delta(1-\gamma)}\right)-2\bigg(c_{2}-c_{3}\bigg(1-\frac{1}{\rho}\bigg)\bigg)\mu\right)$.
Consequently, when all the constants involved are such that $0<\eta<\frac{1}{\bar{\alpha}}$, the convergence rate is linear with an $\mathcal{O}(\log(\frac{1}{\epsilon}))$ iteration, stochastic function evaluations and stochastic gradient evaluations complexity to obtain an $\epsilon$-accurate solution in expectation.
Proof.
The sequence of directions $\{d_{k}\}$ satisfies the conditions (13)-(14) with constants $c_{1}$ and $c_{2}$ as a result of steps 11-14. By Proposition 5.2 and the instructions of the algorithm, it also satisfies condition (15) with $c_{3}$ defined as in the statement of the theorem. All the assumptions of [lapucci2024convergenceconditionsstochasticline, Theorem 5.3] are thus satisfied by the ones here, and we consequently get the thesis.
@@ 6 Computational Experiments
In this section, we present the results of a comprehensive experimental analysis of the proposed Mini-Batch Conjugate Gradient method with Data Persistency (MBCG-DP). As previously discussed in Section 5.1, the bias correction term required to make the gradient estimate unbiased actually opposes to the effect we introduce with the overlap strategy (for completeness, we show in Figure 3 that bias correction indeed slows down the algorithm on the MNIST task considered in Section 3.2). In the following computational experiments we therefore consider the algorithm version without any bias correction. All experiments reported in this section were implemented in Python3333The code for the proposed method is available at github.com/dadoPuccio/MB-Conjugate-Gradient-DP and run on a machine with the following characteristics: Ubuntu 22.04, Intel i5-13400F Processor, 32 GB RAM, Nvidia GeForce RTX 4060 8GB.
Figure 3: Behavior of MBCG-DP and its unbiased variant when employed to train a multi-layer perceptron on MNIST dataset with batch size of 128 and 512.
@@
In Figure LABEL:fig:CG_rule_comparison we reported the behavior of MBCG-DP when comparing the CG rules of Fletcher-Reeves (FR), Polyak-Polak-Ribiére (PPR) and Hestenes-Stiefel (HS). In this experiment the generalized SPS (7) with $c=1$ is used to select $\alpha_{0}^{k}$ and the momentum clipping approach (contraction factor 0.5) from [fan2023] is employed in case $d_{k}$ is not a descent direction. The Fletcher-Reeves rule clearly resulted the most effective one in both tasks and for both batch sizes.
@@
In Figure LABEL:fig:CG_tentative_step_comparison we show the behavior of MBCG-DP when using different approaches for selecting the tentative step $\alpha_{0}^{k}$ for the line search: a constant value, the heuristic employed in [vaswani2019, fan2023] and the generalized SPS (7) with $c=1$ were tested. The Fletcher-Reeves rule is used in this experiment; in case $d_{k}$ is not a descent direction the repeated momentum clipping approach is used, again with contraction factor of $0.5$. The generalized SPS proved to be the most effective choice.
Finally, in Figure LABEL:fig:CG_recover_direction we observe the impact of different recovery options to retrieve a descent direction, namely: switching to the simple (negative) stochastic gradient direction, recomputing the direction according to the subspace optimization strategy proposed in [lapucci2024globally], inverting the direction and taking $d_{k}$ and adopting the repeated momentum clipping approach (contraction factor of $0.5$) from [fan2023]. In this experiment, the Fletcher-Reeves rule is employed in combination with the generalized SPS initial step size. The clipping strategy resulted the preferable choice especially in nonconvex problems.
This detailed comparison allows us to identify a promising setup to run MBCG-DP in further experiments to compare it to state-of-the-art methods. In what follows, we thus employed the Fletcher-Reeves rule for $\beta_{k}$, the generalized stochastic Polyak step size (7) for choosing $\alpha_{0}^{k}$ and the clipping strategy on $\beta_{k}$ as a recovery option for obtaining a descent direction. The resulting method will be denoted by the acronym MBCG_FR.
The other algorithms included in the following experiments are: vanilla SGD with momentum, Adam [kingma2015], SLS [vaswani2019], PoNoS [galli2023] and MSL_SGDM [fan2023]. The benchmark of test problems include three convex instances, consisting in the training of an RBF-kernel classifier on mushrooms, ijcnn and rcv1 datasets444datasets available at csie.ntu.edu.tw/ cjlin/libsvmtools/datasets. Moreover, three nonconvex learning tasks are taken into account: once again a multi-layer perceptron with 1000 hidden units and ReLU activations to be trained on MNIST, a three-layered convolutional architecture with 32 channels each, followed by a fully-connected layer of 512 units to be trained on FashionMNIST [xiao2017fashionMNIST] dataset555available at github.com/zalandoresearch/fashion-mnist and a ResNet18 architecture to be trained on CIFAR10 [krizhevsky2009learning] dataset666available at cs.toronto.edu/ kriz/cifar.html. For the latter network, we employed the normalization-free setup proposed in [civitelli2023]. Batch sizes of 128 and 512 were again considered for all problems.
Table 2: Mini-batch overlap settings leading to the best-performing algorithms across problems.
@@
SGD+M 75% 75% 75% 75% 75% 75% 75% 75% 75% 75% 0% 0%
Adam 75% 75% 75% 75% 75% 75% 50% 50% 75% 75% 0% 10%
SLS 75% 75% 50% 75% 75% 75% 0% 75% 0% 0% 0% 0%
PoNoS 75% 75% 75% 75% 50% 75% 25% 50% 0% 0% 0% 0%
MSL_SGDM 75% 75% 50% 75% 0% 25% 10% 10% 75% 0% 25% 0%
@@
Figure 5: Comparison of the proposed MBCG_FR algorithm with other algorithms in the literature. The training loss over time is shown for each problem.
@@
In order to offer a fair comparison between the proposed MBCG-DP method and its competitors, each reference algorithm was evaluated to assess whether employing mini-batch persistency could enhance its performance. For each problem considered, the reference algorithms are tested using an overlap of 0%, 10%, 25%, 50% and 75%. An algorithm is deemed to perform better with a certain overlap value if it either terminates before the others with a superior solution or, at the termination time of the first configuration to finish, it holds a better solution. The training loss is compared to asses whether a solution is better than the other. We report in Table 2 the outcome of this preliminary analysis. In convex problems, we observe that overlap is almost always beneficial. However, with larger architectures in nonconvex tasks, this benefit does not appear to hold. We argue that, in such cases, the dominant computational burden shifts from data I/O operations to the evaluation of the training loss and its gradient. Since the number of iterations per epoch with a 50% overlap is doubled, in this cases the execution time per epoch will also double. In addition, we remark that hyperparameters were selected for each considered method based on preliminary experiments carried out without employing data persistency. A specialized selection of hyperparameters values for each method and overlap percentage might further enhance performance when data persistency is employed.
@@
Figure 6: Comparison of the proposed MBCG_FR algorithm with other algorithms in the literature. The validation accuracy over time is shown for each problem.